Succinic acid in the Merck Index (hard-copy from Library reserve) or in a MSDS (Material Safety Data Sheet from many Internet all parts
Answer #1

1.

a) Molecular formula of Succinic acid is C4H6O4. It is a dicarboxylic acid (CH₂)₂(CO₂H)₂

b) Its molar mass (in g /mol) is = (4 x 12) + (1 x 6) + (16 x 4) = 118 g/mol

c) Hazard properties of it are: exposure can lead to eye irritation and damage, irritation of skin and respiratory tract. Also, in neutralizing reaction of Succinic acid with organic or inorganic bases, heat energy is produced in a significant amount. The acid is however not toxic or carcinogenic.

2.

If 13.024 g of 5 % solution of succinic acid is used, then it is a weight% measure of the solute i.e. succinic acid.

a) Therefore, amount of succinic acid present in g = (5/100) x (13.024 g) = 0.6512 g

b) and moles = (amount present in g) / (molar mass in g/mol)

                 = (0.6512 / 118 ) = 0.00552